# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ Abstract
The entropic risk measure is widely used in high-stakes decision making to account for tail risks associated with an uncertain loss. With limited data, the empirical entropic risk estimator, i.e. replacing the expectation in the entropic risk measure with a sample average, underestimates the true risk. To debias the empirical entropic risk estimator, we propose a strongly asymptotically consistent bootstrapping procedure. The first step of the procedure involves fitting a distribution to the data, whereas the second step estimates the bias of the empirical entropic risk estimator using bootstrapping, and corrects for it. We show that naively fitting a Gaussian Mixture Model to the data using the maximum likelihood criterion typically leads to an underestimation of the risk. To mitigate this issue, we consider two alternative methods: a more computationally demanding one that fits the distribution of empirical entropic risk, and a simpler one that fits the extreme value distribution. As an application of the approach, we study a distributionally robust entropic risk minimization problem with type-$\infty$ Wasserstein ambiguity set, where debiasing the validation performance using our techniques significantly improves the calibration of the size of the ambiguity set. Furthermore, we propose a distributionally robust optimization model for a well-studied insurance contract design problem. The model considers multiple (potential) policyholders that have dependent risks and the insurer and policyholders use entropic risk measure. We show that cross validation methods can result in significantly higher out-of-sample risk for the insurer if the bias in validation performance is not corrected for. This improvement can be explained from the observation that our methods suggest a higher (and more accurate) premium to homeowners.
@@ 1 Introduction
The purpose of a risk measure is to assign a real number to a random variable, representing the preference of a risk-averse decision maker towards different risky alternatives. For instance, when faced with multiple options, a decision maker might prefer a guaranteed loss of zero over an uncertain option, even if the latter has a strictly negative expected loss. While this behavior can be explained using the mean-variance criterion (Markowitz, 1952), which balances the expected loss and its fluctuations around the mean, the entropic risk measure offers greater flexibility by incorporating higher moments of the loss distribution.
The entropic risk measure is useful in high-stakes decision-making, where rare events and their associated extreme losses are a significant concern. A key advantage of using entropic risk in multi-stage decision-making is the time-consistency of the optimal policies. There has been significant growth in research on exponential utility functions, which appear in the literature under various names, including entropic risk minimization, tilted empirical risk minimization, constant absolute risk aversion, and as special cases of more general shortfall risk measures and optimized certainty equivalent risk measures (Ben-Tal and Teboulle,, 1986). Applications of these concepts are widespread, particularly in finance (Föllmer and Schied,, 2002, 2011; Smith and Chapman,, 2023), portfolio selection (Brandtner et al.,, 2018; Markowitz,, 2014), revenue management (Lim and Shanthikumar,, 2007), economics (Svensson and Werner,, 1993), operations management (Choi and Ruszczyński,, 2011), robotics (Nass et al.,, 2019), statistics (Li et al.,, 2023), reinforcement learning (Fei et al.,, 2021; Hau et al.,, 2023), risk-sensitive control (Howard and Matheson,, 1972), game theory (Saldi et al.,, 2020), and catastrophe insurance pricing (Bernard et al.,, 2020).
@@
$\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}})):=\begin{cases}\frac{1}{\alpha}\log({\mathbb{E}}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))])&\text{ if }\alpha>0,\\ {\mathbb{E}}_{{\mathbb{P}}}[\ell({{\boldsymbol{\xi}}})]&\text{ if }\alpha=0,\end{cases}$ (1)
where the loss $\ell({{\boldsymbol{\xi}}})$ is transformed by the increasing and convex exponential function, and $\alpha$ is the risk aversion parameter. This formulation expresses the entropic risk as the certainty equivalent of the expected disutility $\mathbb{E}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]$, reflecting the monetary value of the risk inherent in the uncertain outcome ${{\boldsymbol{\xi}}}$. By adjusting $\alpha$, the decision maker’s sensitivity to extreme losses can be controlled.
@@
$\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}})):=\frac{1}{\alpha}\log(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))),$ (2)
Since $\log$ function is strongly concave, Jensen’s inequality implies that the empirical entropic risk strictly underestimates the true entropic risk:
${\mathbb{E}}[\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))]\;=\;{\mathbb{E}}[\frac{1}{\alpha}\log(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i})))]<\frac{1}{\alpha}\log({\mathbb{E}}[\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))])=\rho(\ell({{\boldsymbol{\xi}}})),$ (3)
@@
A fundamental challenge in high-stakes decision-making is accurately estimating the risk. Even with a large number of samples, the empirical entropic risk may still significantly underestimate the true risk, particularly for risk-averse decision makers. This issue is illustrated in the following example.
@@
Figure 1: Statistics of the sample risk for different values of the risk aversion parameters $\alpha\in\{0,0.5,1,1.5,2\}$ and training sample sizes $N\in\{100,1000,10000\}$ over 500 repetitions. The true risk is given by $\frac{1}{\alpha}\log\left(\left(1-0.5\alpha\right)^{-10}\right)$
@@
In an insurance pricing problem, the insurer aims to determine the minimum premium $\pi$ at which they can insure against the loss $\ell(\xi):=\xi$. Let risk aversion parameter of insurer be $\alpha$. Assuming full coverage for the losses $\xi$, loss of the insurer if they charge a premium $\pi$ is a random variable given by $\pi-\xi$. Thus, minimum premium at which the insurer insures the risk should be such that the entropic risk of the insurer from insuring is at most equal to 00, i.e., $\frac{1}{\alpha}\log\left({\mathbb{E}}_{\mathbb{P}}[\exp(\alpha(\pi-\xi))]\right)=0$. On rearranging the terms, we can show that the minimum premium equals the entropic risk associated with the loss $\xi$, $\pi=\frac{1}{\alpha}\log\left({\mathbb{E}}_{\mathbb{P}}\left[\exp(\alpha\xi)\right]\right)$. Suppose that the loss follows a $\Gamma(\kappa,\lambda)$ distribution (see Fu and Moncher,, 2004; Bernard et al.,, 2020) with shape parameter $\kappa$ and scale parameter $\lambda$. The moment-generating function of $\Gamma$-distributed random variable is known in closed form which allows us to analytically compute the optimal premium $\pi=\frac{1}{\alpha}\log\left(\left(1-\lambda\alpha\right)^{-\kappa}\right).$
@@
Suppose an insurer has access to $N\in\{100,1000,10000\}$ samples of the losses which are generated from a $\Gamma(10,0.45)$-distribution. We use empirical distribution ^$\hat{{\mathbb{P}}}_{N}$ over $N$ samples to estimate the entropic risk for different risk averse parameters, $\alpha\in\{0,0.5,1,1.5,2\}$. Figure 1 presents statistics of the distribution of the empirical risk estimator as a function of $N$ and $\alpha$. We can see that more risk-averse insurers significantly underestimate the risk of the loss $\xi$ and thus the premium to impose on the insuree. Furthermore, the slow convergence of sample mean to the true mean for heavy tailed random variables is a well-known phenomenon (Catoni,, 2012; Lugosi and Mendelson,, 2019). Even when the underlying distribution of $\ell({{\boldsymbol{\xi}}})$ is light-tailed, $\exp(\alpha\ell({{\boldsymbol{\xi}}}))$ can be heavy-tailed (Nair et al.,, 2022). For instance, $\exp(\alpha\ell({{\boldsymbol{\xi}}}))$ is a lognormally distributed when $\ell({{\boldsymbol{\xi}}})$ is Gaussian random variable.
@@
The contribution of this paper is to propose a scheme to correct for the bias of the empirical entropic risk estimator, that is, find bias correction term $\delta(\mathcal{D}_{N})$ that depends on a distribution fitted to the data $\mathcal{D}_{N}$ so that
@@
$\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))\approx{\mathbb{E}}_{\mathcal{D}_{N}}\left(\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))+\delta(\mathcal{D}_{N})\right).$ (4)
We show that $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))+\delta(\mathcal{D}_{N})\rightarrow\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))$ almost surely with $N\rightarrow\infty$. This suggests that, with an infinite number of samples, it is possible to fit a distribution to the data using MLE and correct the bias. However, with limited number of samples, fitting a distribution using maximum likelihood criteria can lead to estimation errors and we will show later empirically that the underestimation problem persists. Instead, we provide two procedures to learn ”bias-aware” distributions that take into account the bias due to the tail events. The first one involves a distribution matching technique that mimics the bias in the samples, and the second one fits an extreme value distribution to the tails of the data.
@@
Going beyond the estimation of entropic risk, we study the entropic risk minimization problem. Solving the sample average approximation (SAA) of the entropic risk minimization problem is known to produce a second source of bias, also known as the optimizer’s curse (Smith and Winkler,, 2006). Distributionally robust optimization (DRO) is widely used to address the optimistic bias of SAA policies because the decision-maker is protected against perturbations in the empirical distribution that lie in a distributional ambiguity set. Most of the literature on DRO with the Wasserstein ambiguity set assumes that the random variables involved in the expectation operator have light tails, a condition that is not satisfied for the entropic risk measure. It is well-known that worst-case loss in a DRO problem with type-p Wasserstein ambiguity set $(p<\infty)$ is finite if and only if the loss function satisfies a growth condition (Gao and Kleywegt,, 2023). Thus, the worst-case entropic risk can be shown to be unbounded for type-p Wasserstein ambiguity set with $p\in[1,\infty)$. So, we introduce a distributionally robust entropic risk minimization problem with type-$\infty$ Wasserstein ambiguity set and cast it into a finite-dimensional convex optimization problem for piecewise concave loss functions.
To tune the radius of the ambiguity set, a typical approach is to use K-Fold cross validation (CV). We use our bias correction procedure to estimate the validation performance of decisions. This leads to orders of magnitude improvement in out-of-sample entropic risk in terms of optimality gap compared to the “traditional” K-Fold procedure that chooses the radius of distributional ambiguity set based only on evaluating the decision on the validation loss scenarios.
@@
    Our distributionally robust entropic risk minimization problem with a type-$\infty$ Wasserstein ambiguity set has bounded worst-case losses. We obtain its tractable robust counterpart for piecewise concave loss functions using Fenchel duality and provide conditions under which the optimal risk from the the DRO problem converges to the true optimal risk.
@@
    On the application side, our work contributes toward data-driven designing of insurance premium pricing and coverage policies. To the best of our knowledge, this is the first time that a distributionally robust version of the well-known risk-averse insurance pricing problem is introduced in the literature. Our model takes into account the different risk aversion attitudes of the insurer and homeowners as well as the systemic risk associated with catastrophe events.

### v3

# Mitigating optimistic bias in entropic risk estimation and optimization with an application to insurance

@@ Abstract
The entropic risk measure is widely used in high-stakes decision making to account for tail risks associated with an uncertain loss. With limited data, the empirical entropic risk estimator, i.e. replacing the expectation in the entropic risk measure with a sample average, underestimates the true risk. To mitigate the bias in the empirical entropic risk estimator, we propose a strongly asymptotically consistent bootstrapping procedure. The first step of the procedure involves fitting a distribution to the data, whereas the second step estimates the bias of the empirical entropic risk estimator using bootstrapping, and corrects for it. Two methods are proposed to fit a Gaussian Mixture Model to the data, a computationally intensive one that fits the distribution of empirical entropic risk, and a simpler one with a component that fits the tail of the empirical distribution. As an application of our approach, we study distributionally robust entropic risk minimization problems with type-$\infty$ Wasserstein ambiguity set and apply our bias correction to debias validation performance. Furthermore, we propose a distributionally robust optimization model for an insurance contract design problem that takes into account the correlations of losses across households. We show that choosing regularization parameters based on the cross validation methods can result in significantly higher out-of-sample risk for the insurer if the bias in validation performance is not corrected for. This improvement in performance can be explained from the observation that our methods suggest a higher (and more accurate) premium to homeowners.
@@ 1 Introduction
The purpose of a risk measure is to assign a real number to a random variable, representing the preference of a risk-averse decision maker towards different risky alternatives. For instance, when faced with multiple options, a decision maker might prefer a guaranteed loss of zero over an uncertain option, even if the latter has a strictly negative expected loss. While this behavior can be explained using the mean-variance criterion (Markowitz,, 1952), which balances the expected loss and its fluctuations around the mean, the entropic risk measure offers greater flexibility by incorporating higher moments of the loss distribution.
The entropic risk measure is useful in high-stakes decision-making, where rare events and their associated extreme losses are a significant concern. A key advantage of using entropic risk in multi-stage decision-making is the time-consistency of the optimal policies. The entropic risk measure is widely used due to its interpretation as the certainty equivalent of the exponential utility function (Von Neumann and Morgenstern,, 1944), which represents the risk preferences of a decision-maker exhibiting constant absolute risk aversion (CARA – Arrow,, 1971; Pratt,, 1964). There has been significant growth in research on exponential utility functions, which appear in the literature under various names, including entropic risk minimization, tilted empirical risk minimization, constant absolute risk aversion, and as special cases of more general shortfall risk measures and optimized certainty equivalent risk measures (Ben-Tal and Teboulle,, 1986). Applications of these concepts are widespread, particularly in finance (Föllmer and Schied,, 2002, 2016; Smith and Chapman,, 2023), portfolio selection (Brandtner et al.,, 2018; Markowitz,, 2014; Chen et al., 2024b, ), revenue management (Lim and Shanthikumar,, 2007), economics (Svensson and Werner,, 1993), operations management (Choi and Ruszczyński,, 2011; Chen and Sim,, 2024), robotics (Nass et al.,, 2019), statistics (Li et al.,, 2023), reinforcement learning (Fei et al.,, 2021; Hau et al.,, 2023), risk-sensitive control (Howard and Matheson,, 1972; Bäuerle and Jaśkiewicz,, 2024), game theory (Saldi et al.,, 2020), and catastrophe insurance pricing (Bernard et al.,, 2020).
@@
$\rho^{\alpha}_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}})):=\begin{cases}\frac{1}{\alpha}\log({\mathbb{E}}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))])&\text{ if }\alpha>0,\\ {\mathbb{E}}_{{\mathbb{P}}}[\ell({{\boldsymbol{\xi}}})]&\text{ if }\alpha=0,\end{cases}$ (1)
where the loss $\ell({{\boldsymbol{\xi}}})$ is transformed by the increasing and convex exponential function, and $\alpha$ is the risk aversion parameter. This formulation expresses the entropic risk as the certainty equivalent of the expected disutility $\mathbb{E}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]$, reflecting the monetary value of the risk inherent in the uncertain outcome $\ell({{\boldsymbol{\xi}}})$. By adjusting the risk-aversion parameter $\alpha$, also known as the Arrow-Pratt measure of risk aversion, the decision maker’s sensitivity to extreme losses can be controlled. For the remainder of the paper, we simplify the notation by suppressing the dependence on $\alpha$ where it is not essential, and denote entropic risk as $\rho_{\mathbb{P}}$.
@@
$\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}})):=\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))\right).$ (2)
Since the logarithm function is strongly concave, Jensen’s inequality implies that the empirical entropic risk strictly underestimates the true entropic risk:
${\mathbb{E}}[\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))]\;=\;{\mathbb{E}}\left[\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))\right)\right]<\frac{1}{\alpha}\log\left({\mathbb{E}}\left[\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))\right]\right)=\rho(\ell({{\boldsymbol{\xi}}})),$ (3)
@@
A fundamental challenge in high-stakes decision-making lies in accurately estimating risk. Even with a large number of samples, the empirical entropic risk can significantly underestimate the true risk, especially for decision makers with heightened risk sensitivity. This challenge is demonstrated in the following example.
@@
Figure 1: Statistics of the empirical risk for different values of the risk aversion parameters $\alpha\in\{0,0.5,1,1.5,2\}$ and training sample sizes $N\in\{50,100,200,500\}$ over $10000$ repetitions. The true risk is given by $\frac{1}{\alpha}\log\left(\left(1-0.24\alpha\right)^{-10}\right)$.
@@
In an insurance pricing problem, the insurer aims to determine the minimum premium $\pi$ at which they can insure against the loss $\ell(\xi):=\xi$. Let risk aversion parameter of insurer be $\alpha$. Assuming full coverage for the losses $\xi$, loss of the insurer if they charge a premium $\pi$ is a random variable given by $\pi-\xi$. Thus, minimum premium at which the insurer insures the risk should be such that the entropic risk of the insurer from insuring is at most equal to 00, i.e., $\frac{1}{\alpha}\log\left({\mathbb{E}}_{\mathbb{P}}[\exp(\alpha(\pi-\xi))]\right)=0$. On rearranging the terms, one can show that the minimum premium equals the entropic risk associated with the loss $\xi$, $\pi=\frac{1}{\alpha}\log\left({\mathbb{E}}_{\mathbb{P}}\left[\exp(\alpha\xi)\right]\right)$, also called the exponential premium (Gerber,, 1974). Suppose that the loss follows a Gamma distribution $\Gamma(\kappa,\lambda)$ (see Fu and Moncher,, 2004; Bernard et al.,, 2020) with shape parameter $\kappa$ and scale parameter $\lambda$. The moment-generating function of $\Gamma$-distributed random variable is known in closed form which allows us to analytically compute the optimal premium $\pi=\frac{1}{\alpha}\log\left(\left(1-\lambda\alpha\right)^{-\kappa}\right)$ if $\lambda<1/\alpha$.
@@
Suppose an insurer has access to $N\in\{50,100,200,500\}$ samples of the losses which are generated from a $\Gamma(10,0.24)$-distribution. We use empirical distribution ^$\hat{{\mathbb{P}}}_{N}$ over $N$ samples to estimate the entropic risk for different risk aversion parameters, $\alpha\in\{0,0.5,1,1.5,2\}$. Figure 1 presents statistics of the distribution of the empirical risk estimator as a function of $N$ and $\alpha$. We can see that more risk-averse insurers significantly underestimate the risk of the loss $\xi$ and thus the premium to impose on the insuree even for the relatively large sample size $N=500$. This phenomenon can also be understood by analyzing the impact of an infinitesimal change in the distribution on the entropic risk, as quantified by the influence function (Hampel,, 1974). For this example, the influence function is given by $\text{IF}(\hat{\xi})=-\frac{1}{\alpha}+\frac{\exp(\alpha\hat{\xi})}{\alpha\left(1-0.24\alpha\right)^{-10}}$. Consequently, missing tail scenarios can disproportionately affect the estimation of entropic risk for high values of $\alpha$ due to the exponential term. Detailed calculations are provided in Appendix B.1, and Figure 14 offers a visual representation of the influence function across samples and their associated probabilities. The slow convergence of sample mean to the true mean for heavy tailed random variables is a well-known phenomenon (Catoni,, 2012; Lugosi and Mendelson,, 2019). Even when the underlying distribution of $\ell({{\boldsymbol{\xi}}})$ is light-tailed, $\exp(\alpha\ell({{\boldsymbol{\xi}}}))$ can be heavy-tailed (Nair et al.,, 2022). For instance, $\exp(\alpha\ell({{\boldsymbol{\xi}}}))$ is a lognormally distributed when $\ell({{\boldsymbol{\xi}}})$ is Gaussian random variable.
@@
The contribution of this paper is to propose a scheme to produce (approximately) median-unbiased estimators of entropic risk to mitigate the bias of the empirical entropic risk estimator. We propose a bias correction term $\delta(\mathcal{D}_{N})$ that employs bootstrapping, using a distribution fitted to the data $\mathcal{D}_{N}$, to get
@@
$\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))\approx\text{median}\left[\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))+\delta(\mathcal{D}_{N})\right],$ (4)
where the median statistic is taken with respect to the randomness in $\mathcal{D}_{N}$. We establish mild conditions under which $\delta(\mathcal{D}_{N})\rightarrow 0$ almost surely as $N\rightarrow\infty$. In particular, basing the bootstrap on a constrained maximum likelihood estimation (MLE) of sampling distribution will yield an asymptotically consistent estimator. Unfortunately, our empirical experiments establish that a bootstrap correction based on MLE fails to adequately address the underestimation of entropic risk. Instead, we provide two procedures to fit “bias-aware” distributions that take into account the entropic risk estimation bias caused by tail events. The first one involves a distribution matching technique that tries to fit the entropic risk estimator’s distribution itself, and the second uses a simple mixture distribution with a component dedicated to fitting the tail of the empirical distribution.
@@
Going beyond the estimation of entropic risk, we study the entropic risk minimization problem. Solving the sample average approximation (SAA) of the entropic risk minimization problem is known to produce a second source of bias, also known as the optimizer’s curse (Smith and Winkler,, 2006). Distributionally robust optimization (DRO) is widely used to address the optimistic bias of SAA policies because the decision-maker is protected against perturbations in the empirical distribution that lie in a distributional ambiguity set. Most of the literature on DRO with the Wasserstein ambiguity set assumes that the random variables involved in the expectation operator have light tails, a condition that is not satisfied for the entropic risk measure. It is well-known that worst-case loss in a DRO problem with type-p Wasserstein ambiguity set $(p<\infty)$ is finite if and only if the loss function satisfies a growth condition (Gao and Kleywegt,, 2023). Thus, the worst-case entropic risk can be shown to be unbounded for type-p Wasserstein ambiguity set with $p\in[1,\infty)$. So, we introduce a distributionally robust entropic risk minimization problem with type-$\infty$ Wasserstein ambiguity set and cast it into a finite-dimensional convex optimization problem for piecewise concave loss functions. To tune the radius of the ambiguity set, a typical approach is to use K-Fold cross validation (CV). We use our bias mitigation procedure to estimate the validation performance of the resulting decisions.
To demonstrate the effectiveness of our approach, we conduct a case study on an insurance pricing problem. The International Panel on Climate Change (IPCC) advocates using financial instruments like catastrophe insurance to mitigate risks associated with rare, high-impact events such as floods, earthquakes, and wildfires. These events have become more frequent due to climate change, making reliable risk estimation even more critical (Linnerooth-Bayer and Hochrainer-Stigler,, 2015). The correlated risks of floods, earthquakes, and wildfires can lead to significant payouts for insurers, limiting the availability of insurance for such events in commercial markets (Marcoux and H Wagner,, 2023). The true distribution of losses, which may be correlated across homeowners, is unknown to both the insurer and the homeowners. The insurer addresses this uncertainty by solving a distributionally robust insurance pricing problem, determining the coverage to offer and the premium to charge each homeowner. Our results show that the insurer can achieve a significant improvement in out-of-sample entropic risk compared to the “traditional” K-Fold CV procedure, which selects the radius of the distributional ambiguity set solely by evaluating decisions on validation loss scenarios.
@@
    We introduce a distributionally robust entropic risk minimization problem with a type-$\infty$ Wasserstein ambiguity set with bounded worst-case losses. We obtain its tractable robust counterpart for piecewise concave loss functions using Fenchel duality and provide conditions under which the optimal risk from the the DRO problem converges to the true optimal risk.
@@
    On the application side, our work contributes toward data-driven designing of insurance premium pricing and coverage policies. To the best of our knowledge, this is the first time that a distributionally robust version of the well-known risk-averse insurance pricing problem (Bernard et al.,, 2020) is introduced in the literature. Our model takes into account the different risk aversion attitudes of the insurer and homeowners as well as the systemic risk associated with catastrophe events.